put this conic section in standard form. What is the a^2 value? 25x^2+16y^2+150x=160y-225
Answer by lwsshak3(6519)

put this conic section in standard form. What is the a^2 value? 25x^2+16y^2+150x=160y-225
**
25x^2+16y^2+150x=160y-225
25x^2+150x+16y^2-160y=225
complete the squares
25(x^2+6x+9)+16(y^2-10+25)=-225+225+400
(x+3)^2/16+(y-5)^2/25=1
This is an equation for an ellipse with vertical major axis of the standard form:
(x-h)^2/b^2+(y-k)^2/a^2=1, a>b, (h,k) being the (x,y) coordinates of the center.
For given equation:
a^2=25
a=5